OpenStudy (anonymous):

Finding combinations Evaluate each expression Ten candidates are running for three seats in the student government. You may vote for as many as three candidates. In how many ways can you vote

OpenStudy (anonymous):

\[\binom{10}{0}+\binom{10}{1}+\binom{10}{2}+\binom{10}{3}\]

OpenStudy (anonymous):

\[1+10+\frac{10\times 9}{2}+\frac{10\times 9\times 8}{3\times 2}\]

OpenStudy (anonymous):

i dont understand what meant \[\( 10 \0) \] it is when he dont vote for any candidate

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but it must no count

OpenStudy (anonymous):

oh! so one way that he can vote is for any of them

OpenStudy (anonymous):

one way he can vote for no one ten ways to vote for one etc

OpenStudy (anonymous):

i got thanks